OpenStudy (lexy148):

convert mole to gram problems 1.62moles of Fe2(SO3)3 & 1.35moles of Mg(NO3)2

OpenStudy (anonymous):

find the molar mass of Fe2(so3)3 first, then use the molar mass times 1.62

OpenStudy (wolfe8):

For both, use n=m/molar mass

OpenStudy (lexy148):

would it be Fe=2(55.847) + s=3(32.066) + 3

OpenStudy (wolfe8):

What about O?

OpenStudy (lexy148):

o=15.999

OpenStudy (lexy148):

so how would i add up all those numbers i need help

OpenStudy (wolfe8):

So the molar mass of the first compound will be 2(55.84)+3(32.08+3(15.99))

OpenStudy (lexy148):

oh okay thx

OpenStudy (wolfe8):

You're welcome. Can you do this now?

OpenStudy (lexy148):

yes would you do the same process for the 2nd problem

OpenStudy (wolfe8):

Yup but of course with the respective atomic masses. Also, this only gives you the molar mass. You want the mass. You know this, right?

OpenStudy (lexy148):

yes i know how to do the rest thank you you helped alot

OpenStudy (wolfe8):

No problem. :)

OpenStudy (lexy148):

one last question would it equal 351.83 for the sum of the elements shown

OpenStudy (wolfe8):

Yep :)

OpenStudy (lexy148):

k just making sure

OpenStudy (wolfe8):

Yup no problem.

OpenStudy (wolfe8):

Do take time to read our CoC: http://openstudy.com/code-of-conduct and enjoy your stay. If you are satisfied with any question, please reward a medal to whoever you think deserves it and close the question. We also have a fan and testimony system. See you around :)